11-722: Grammar Formalisms

Spring Term 2008 Language Technologies Institute

Carnegie Mellon University

Course Syllabus

Time and Place:

T-Th, 3:00-4:30  Wean Hall 4623

Instructors:

Lori Levin  Associate Research Professor

Language Technologies Institute Newell-Simon Hall 4631 412-268-6193

lsl@cs.cmu.edu

Alan Black  Associate Research Professor

Language Technologies Institute 4525 Newell-Simon Hall 412-268-6299

awb@cs.cmu.edu

Teaching Assistant

Rashmi Gangadharaiah Ph.D. Student

Language Technologies Institute

rgangadh@andrew.cmu.edu

Prerequisites:

Permission of instructors, or background in linguistics, computational linguistics, language technologies, or psycholinguistics.

Course Description:

The goal of this course is to familiarize students with grammar formalisms (and the underlying linguistic theories) that are commonly used for research in computational lingusitics, language technologies, and lingusitics. We hope to have students from a variety disciplines (linguistics, computer science, psychology, modern languages, philosophy) in order to cover a broad perspective in class discussions. Comparison of formalisms and the underlying theories will lead to a deeper understanding of human language and natural language processing algorithms. The formalisms will include: Head Driven Phrase Structure Grammar, Lexical Functional Grammar, Tree Adjoining Grammar, and Categorial Grammar. We will also cover Penn Treebank, Prague Treebank, CCG treebank, and other treebanks as time permits. We will cover the treatment of basic syntactic and semantic phenomena in each formalism, and will focus on how to recognize and analyze these phenomena in real text (as opposed to artificial text-book examples). We will also cover grammar induction from treebanks.

Requirements: Four problem sets (40%), a term project (50%), and class attendance and participation (10%).

Texts:

• Lexical Functional Syntax, Joan Bresnan, Blackwell, 2000.

• Lexical-Functional Grammar (Syntax and Semantics, Volume 34) by Mary Dalrymple, Academic Press.

• Syntactic Theory: A Formal Introduction by Ivan A. Sag and Thomas Wasow (Author), CSLI, 1999.

• Other texts will be added.

Schedule: Part One: Introduction

Jan. 15 Review syllabus and schedule  Black & Levin
Jan. 17 History of grammar formalisms  Levin
Jan. 22 Syntax Checklist and Penn Treebank Levin
Jan. 24 Checklist and Treebank  Levin
Jan. 29 Checklist and Treebank  Levin
Jan. 31 Checklist and Treebank  Levin
Feb. 5 Checklist and Treebank  Levin

Part Two: The Formalisms

Feb. 7  Categorial Grammar  Black

Formalism and Lambda Calculus

Feb. 12 Categorial Gammar Black

Syntax Checklist

Feb. 14 Categorial Grammar Black

Syntax Checklist

Feb. 19 Categorial Grammar Black

Grammar Induction

Feb. 21 Categorial Grammar Black

Grammar Induction

Feb. 26  HPSG  Levin
Feb. 28  HPSG  Levin
Mar. 4  HPSG  Levin
Mar. 6  HPSG  Levin
Mar. 11  Spring Break
Mar. 13  Spring Break
Mar. 18  HPSG  Levin
Mar. 20  HPSG  Levin
Mar. 25  LFG  Levin
Mar. 27  LFG  Levin
Apr. 1  LFG  Levin
Apr. 3  LFG  Levin
Apr. 8  LFG  Levin
Apr. 10  LFG  Levin
After Apr. 15 To be announced